# Modelling the interaction between ethnicity and infectious disease transmission dynamics

### v1

@@ 2.1.1 Proportionate mixing
We first explored the simplest case that assumed each person’s contacts were distributed across ethnicity groups in proportion to their population size. This defined each element of the per capita transmission matrix as proportional to the transmission rates of two ethnicities multiplied together, as seen in Equation 10. This left four parameters to fit, the values of $a_{k}$.
@@
The population of ethnic group $k$ is represented with $N_{k}$.
@@ 2.1.2 Assortative mixing
We then considered a more realistic representation of transmission by accounting for preferential interaction within ethnicities with the introduction of the assortativity coefficient, $\epsilon\in[0,1]$, which was the same for all ethnicities. $\epsilon=0$ is proportionate mixing, $\epsilon=1$ is purely within group, and intermediate values of $\epsilon$ interpolate between these two extremes. This helped to model community and family interactions that were primarily intra-ethnic interactions. We then defined the assortative per capita transmission matrix via Equation 11.
@@
$\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise. As can be seen, swapping i and j does not alter the equation, which means that the per capita transmission matrix was symmetric. From this equation, it can be seen that proportionate mixing was a special case where $\epsilon=0$.
@@ 2.1.3 Initialisation
To initialise the model, we assumed that 0.01% of each ethnicity was exposed to the infection, which gave approximately 500 initially exposed people.
@@
$I_{k}(0)$ $=R_{k}(0)=0$
$N_{vk}$ refers to the vaccinated population of ethnic group $k$, and $w=0.0001$ was the initial proportion of the population exposed to the disease.
@@ 2.2.1 COVID-19 case data
The data for the number of confirmed COVID-19 cases was downloaded from the Populations web tool (30 April, 2025), Health New Zealand, based on customised population projections provided by Stats NZ according to assumptions agreed to by Health New Zealand [30]. This data grouped the number of new cases per ethnicity for each day using a priority measure to assign each case to a single ethnicity (Figure 1), which meant in the case of individuals belonging to more than one ethnic group, their case was assigned in the priority order of Māori, Pacific Peoples, Asian, and finally European/Other. The bulk of confirmed cases during the first Omicron wave were self-reported RATs (Rapid Antigen Tests), with some PCR (Polymerase Chain Reaction) tests (for cases presenting to healthcare). We assumed values for the case ascertainment rate (CAR), the proportion of cases that were reported, to account for asymptomatic cases and unreported cases (as the bulk of cases were self reported). As the true values for CAR were not known, we assumed four different situations: equal CARs of 40%, 50%, and 60% across ethnicities, and a situation where Māori and Pacific Peoples reported fewer cases and had a 40% CAR, while Asian and Europeans/Others had a CAR of 60%. We included this scenario as it is possible that Māori and Pacific Peoples had lower testing rates due to factors such as reduced access to healthcare services, higher rates of insecure employment, financial pressure, and/or inability to work remotely. CAR estimates were based on an Aotearoa study from February-July 2022 using wastewater surveillance. In this study, 40% of 20–25-year-olds reported a case, giving an approximate lower bound on CAR, and model comparison to border workers gave an upper bound of around 60% [32]. It can be seen in Figure 1 that there is a sharp increase in the reported Pacific cases present when self-testing became more available. This could imply that a significant portion of early Pacific cases could not be reported due to limited test availability and are therefore missing from the dataset.
@@ 2.2.2 Vaccinated population data
The Ministry of Health GitHub repository [15] was used to obtain values for the vaccinated populations for each ethnicity. The dataset contained the number of first doses administered in Aotearoa by the 16th of February, 2022, which was used as the vaccinated population. The proportion of coverage was similar among 1st and 2nd dose numbers, being 57.9%, 76.0%, 75.3%, and 84.4% compared to 55.3%, 74.4%, 74.6%, and 83.4% for Māori, Pacific Peoples, Asians, and Europeans/Other, respectively.
2.2.3 Proportionate mixing estimation
Ethnicity-specific attack rates (i.e., cumulative number of cases per capita) were calculated from the 1st of February to the 16th of May, 2022, to capture the first Omicron wave in Aotearoa. These attack rates were then scaled up using the CAR. The four ethnicity-specific transmission rates, $a_{k}$, were estimated by fitting the model to the attack rate data under the proportionate mixing assumption such that the attack rates from the model matched the attack rates from the case data by using the optimize.fsolve function from the SciPy package in Python with its default parameters. All models were fit over a 365-day period, as it allowed sufficient time for the epidemic to run its course.
@@ 2.2.4 Area-level population data
Population data for the populations of each ethnic group in specified geographical areas was needed to fit the assortative model. We used ethnicity-stratified population data at two levels of spatial resolution: Statistical Area 1 (SA1) and Statistical Area 2 (SA2). We assumed these areas represented the bulk of interactions for an individual living within that area.
The SA1 regions were formed by joining meshblocks (the smallest administrative unit) to "allow the release of more detailed information about population characteristics…" [33, p. 13]. The SA1 regions have an ideal population range of 100 to 200 people in each region.
The SA2 regions were defined in a way that "aims to reflect communities that interact together socially and economically" [33, p. 14]. SA2 regions have between 1,000 and 4,000 people in them, but in some cases have less than 1,000. As the SA2 blocks were formed to model a community’s social interactions, they may appear a more accurate measure. However, a substantial amount of COVID-19 infections occurred within households [34], and as such, the finer-scale grouping of the population that gave more weight to local ethnic distributions may give valuable insights. Statistics for the statistical area datasets used can be seen in Appendix A.
The national ethnic population estimates were found by summing over Te Whatu Ora’s District Health Board SA2 populations by prioritised ethnicity [31]. This gave population estimates of 895,040, 363,330, 877,240, and 3,025,980 for Māori, Pacific Peoples, Asians, and Europeans/Others, respectively.
@@
Table of parameter values. Symbol Meaning Value $\boldsymbol{\beta}$ Per capita transmission matrix Method dependent $\sigma$ Inverse of latent period 1/3 day-1 $\gamma$ Inverse of infectious period 1/4 day-1 cumulative reported cases per capita for ethnic group $k$ Māori 23.9% Pacific 30.5% Asian 17.1% European/Other 18.6% $N_{k}$ Population of ethnic group $k$ Māori 888840 Pacific 359480 Asian 820580 European/Other 3048470 $N_{vk}$ Vaccinated population of ethnic group $k$ Māori 514930 Pacific 274413 Asian 629841 European/Other 2558007 $v_{e}$ Vaccine effectiveness 0.3
2.2.5 Assortative mixing estimation
@@
Statistical area data was used to select a value for the assortativity constant using a similar method as Ma et al. [23]. The method first constructed an empirically estimated transmission matrix, $\boldsymbol{T}^{\prime}$, using a specified transmission vector, $\boldsymbol{a}$, as defined in the following equation:
$T_{ij}^{\prime}$ $=\sum_{l}^{L}\frac{a_{i}a_{j}N_{i,l}N_{j,l}}{\sum_{k}a_{k}N_{k,l}}$ (12)
@@
To construct this matrix, a proportionate mixing scheme was applied to the populations of each statistical area to calculate the transmission matrix within that statistical area. The difference between Ma et al.’s method and ours was we assumed frequency-dependent interactions within statistical areas, whereas they assumed density-based interactions in their block groups. Frequency-dependent interactions assume that the mean number of people an individual interacts with each day is constant and does not depend on the population size in the census block or statistical area, while the mean number of density-dependent interactions does. The aggregation over the statistical areas gave the empirically estimated transmission matrix. This was based on an approximating assumption that people mix proportionately with other people in the same geographical area and do not mix with people in other areas. This assumption was not entirely accurate to the real world but gave a good idea of the interactions an individual from each statistical area would have.
@@
$\boldsymbol{a}$ and $\epsilon$ were estimated jointly via an iterative method as follows: For a given $\boldsymbol{a}$, we found the value of $\epsilon$ that minimised the difference between the empirical matrix T’ and the assortative mixing matrix T, defined by the following equation:
$T_{ij}=(1-\epsilon)\frac{a_{i}a_{j}N_{i}N_{j}}{\sum_{k}a_{k}N_{k}}+\epsilon\delta_{ij}a_{i}N_{j}$ (13)
@@
where $N_{i,l}$ is the population of ethnic group $i$ living in geographic region $l$. Then, for this value of epsilon, we estimated $\boldsymbol{a}$ by fitting the model to the attack rate data in the same way as described previously. This iterative process was repeated until the absolute difference between successive values of $\epsilon$ was less than $10^{-6}$.
2.2.6 Non-parametric mixing estimation
The assortative mixing model allowed for within-group contacts to occur more frequently than under the proportionate mixing model but still assumed that the level of assortativity was the same for all ethnic groups and that inter-ethnic contacts occurred in proportion to population size. To investigate the effects of relaxing this assumption, we directly estimated the transmission matrix using the 4 transmission rates and the statistical area data. We did this by constructing the empirically estimated transmission matrix via Equation 12 and using that to fit a model to the attack rates. This meant only four parameters needed to be estimated, the four transmission rates, leaving one unique fitted solution. The fitted empirically estimated transmission matrix was referred to as the non-parametric transmission matrix due to its non-parametric nature. The use of this matrix was made possible by the change to frequency-based interactions within the SAs.
To conserve the transmission rate definition and symmetry in the per capita matrix, the non-parametric matrix required that the populations used to construct it equal the populations used in the model. This meant the non-parametric matrix was only constructed using the prioritised SA2 data.
@@ 2.3 Scenario analysis
We created four scenarios to quantify the effect that the disparity in vaccination, disparity in transmission rates, and the assortativity constant had on model results. These scenarios, seen in Table 2.3, took the assortative model fit and did one of three things: set $\epsilon$ to zero to measure the effect of assortativity on the model; set all transmission rates to the national transmission rates; or set all vaccination rates to the national vaccination rates. These scenarios conserve the total number of vaccinations and the national transmission rates.
@@ 3 Results
Code to replicate results can be found in a GitHub repository [35].
3.1 Proportionate mixing
@@
Figure 2 shows the time series of the susceptible, exposed, infectious, and recovered populations for each ethnicity. As expected with an SEIR model, a single epidemic wave that ends as a result of herd immunity in the population was observed. Recall that the transmission rates were fit to match the attack rates observed in the data. As such, the final recovered population in each ethnicity matched the number of people infected according to the attack rate data and the assumed value for CAR (indicated by the dashed horizontal line).
@@
Figure 2: Per capita SEIR plots fit to data on confirmed cases under the assumption of proportionate mixing and an assumed CAR of 50% for a) Māori, b) Pacific Peoples, c) Asians, and d) Europeans/Others. The dashed line is the cumulative attack rate according to the data on reported cases and the assumed CAR of 50%.
@@
For all the assumed values of CAR, we fit transmission rates to match the new attack values; see Figure 3 and Table 3.2.2. For all values of CAR investigated, Pacific Peoples had the highest transmission rate, followed by Māori, Europeans/Others, and then Asians. The initial reproduction number increased with a decrease in the assumed CAR and was highest when Māori and Pacific Peoples were given a higher CAR than the other groups.
Figure 3: a) Comparison between the transmission rate fits such that the attack rates match the data on confirmed cases for all the assumed CARs under the assumption of proportionate mixing and b) the initial reproduction number for these transmission rates.
3.2 Assortative mixing
3.2.1 Specified $\epsilon$

### v2

@@ 3.1.1 Proportionate mixing
We first explored the simplest case that assumed each person’s contacts were distributed across ethnicity groups in proportion to their population size. This defined each element of the per capita transmission matrix $\beta_{ij}$ as proportional to the transmission rates of two ethnicities multiplied together:
@@
where $N_{k}$ is the population of ethnic group $k$. This left four parameters to estimate, the values of $a_{k}$.
@@ 3.1.2 Assortative mixing
We then considered a more realistic representation of transmission by accounting for preferential interaction within ethnicities with the introduction of the assortativity coefficient, $\epsilon\in[0,1]$, which was assumed to be the same for all ethnicities. When $\epsilon=0$ this corresponds to proportionate mixing, while $\epsilon=1$ is purely within group mixing, and intermediate values of $\epsilon$ interpolate between these two extremes. This helped to model community and family interactions that were primarily intra-ethnic interactions. We defined the assortative per capita transmission matrix as:
@@
where $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise. As can be seen, interchanging $i$ and $j$ did not alter the equation, which meant that the per capita transmission matrix was symmetric.
@@ 3.1.3 Initialisation
To initialise the model, we assumed that a fixed proportion $w$ of each ethnicity was exposed to the infection, with the remaining population in one of the susceptible compartments, giving the following initial conditions:
@@
$E_{vk}(0)$ $=E_{vbk}(0)=I_{k}(0)=I_{vk}(0)=I_{vbk}(0)=R_{k}(0)=0$
$N_{vk}$ refers to the vaccinated population of ethnic group $k$, and $w=0.0001$ was the initial proportion of the population exposed to the disease (giving approximately 500 initially exposed people).
@@ 3.2.1 COVID-19 case data
The data for the number of confirmed COVID-19 cases [te_whatu_ora_covid-19_2025]. This data grouped the number of new cases per ethnicity for each day using a priority measure to assign each case to a single ethnicity (Figure 1), which meant in the case of individuals belonging to more than one ethnic group, their case was assigned in the priority order of Māori, Pacific, Asian, and finally European/Other (values presented in Table 3.2.3). This was opposed to a total measure that would count an individual’s case towards all ethnic groups they identified with. The bulk of confirmed cases during the first Omicron wave were self-reported rapid antigen tests (RATs), with some polymerase chain reaction (PCR) tests (for cases presenting to healthcare). We assumed values for the case ascertainment rate (CAR), the proportion of cases that were reported, to account for asymptomatic cases and unreported cases (as the bulk of cases were self reported). As the true values for CAR were not known, we assumed four different situations: equal CARs of 40%, 50%, and 60% across ethnicities, and a situation where the Māori and Pacific groups reported fewer cases and had a 40% CAR, while the Asian and European/Other groups had a CAR of 60%. We included this scenario as it is possible that Māori and Pacific Peoples had lower testing rates due to factors such as reduced access to healthcare services, higher rates of insecure employment, financial pressure, and/or inability to work remotely. CAR estimates were based on an Aotearoa study from February-July 2022 using wastewater surveillance. In this study, 40% of 20–25-year-olds reported a case, giving an approximate lower bound on CAR, and model comparison to border workers gave an upper bound of around 60% [watson_jointly_2024]. It can be seen in Figure 1 that there was a sharp increase in the reported Pacific cases present when self-testing became available in late February 2022, which followed a period in which PCR testing capacity became overwhelmed [allen__clarke_covid-19_2022]. This could imply that a significant portion of early Pacific cases could not be reported due to limited test availability and are therefore missing from the dataset.
@@ 3.2.2 Vaccinated population data
Ministry of Health data [ministry_of_health_new_2023] was used to obtain values for the vaccinated populations for each ethnicity on the 23rd of February 2022 (see Table 3.2.3). For simplicity, we placed people who had received one or two vaccine doses in the $S_{v}$ compartment, and those who had received a third dose in the $S_{vb}$ compartment. Note that the number of people who had received only one dose was relatively small (1.2%) as most people completed the primary course of two doses.
@@ 3.2.3 Area-level population data
Data for the population size of each ethnic group in specified geographical areas was needed to parameterise the assortative model. We used ethnicity-stratified population data at two levels of spatial resolution: Statistical Area 1 (SA1) and Statistical Area 2 (SA2) level. We assumed that the statistical area an individual lived in contained the bulk of interactions for that individual. StatsNZ reports population numbers at SA1 and SA2 level, which have typical populations of 100-200 and 1,000-4,000 people, respectively. The SA1 regions were formed by joining meshblocks (the smallest administrative unit) to “allow the release of more detailed information about population characteristics…” [statsnz_statistical_2023, p. 13]. The SA2 regions were defined in a way that “aims to reflect communities that interact together socially and economically” [statsnz_statistical_2023, p. 14]. As the SA2 blocks were formed to model a community’s social interactions, they may appear a more accurate measure. However, a substantial amount of COVID-19 infections occurred within households [chen_fangcang_2020], and as such, the finer-scale grouping of the population gave more weight to local ethnic distributions and may give valuable insights. We used two datasets for SA2 population data with different population measures. The first was a dataset containing SA2 data by prioritised-ethnicity that assigned each individual to one ethnicity using the same priority as the COVID-19 case dataset. This data was downloaded from the Populations web tool (30 April, 2025), Health New Zealand, based on customised population projections provided by Stats NZ according to assumptions agreed to by Health New Zealand [te_whatu_ora_populations_2023]; the other was a dataset from Stats NZ with a total measure available at both SA1- and SA2-level that counted an individual towards all ethnic groups that person identified with [statsnz_statistical_2020]. Statistics about each dataset are present in Supplementary Table 1. Maps of the ethnic composition of each SA2 region are presented in Supplementary Figures 2 and 3. The Asian and Pacific populations are clustered primarily in Auckland, which is made up of small, densely population statistical areas, and as such the population appears much lesser than it is. These maps also show the high variability of the geographical areas of SA2s, which are mostly divided according to population.
The national ethnic population estimates (seen in Table 3.2.3) were found by summing over the SA2 populations by prioritised ethnicity.
@@
Table of parameter values. Symbol Meaning Value $\boldsymbol{\beta}$ Per capita transmission matrix Method dependent $\sigma$ Inverse of latent period 1/3 day-1 $\gamma$ Inverse of infectious period 1/4 day-1 $v_{a}$ Complement of vaccine acquisition prevention effectiveness 0.8 $v_{t}$ Complement of transmission prevention effectiveness 0.95 $v_{a}b$ Complement of vaccine and booster acquisition prevention effectiveness 0.55 $v_{t}b$ Complement of vaccine and booster transmission prevention effectiveness 0.85
[…]
Table of population characteristics Symbol Meaning Value Māori Pacific Asian European/Other $N_{k}$ Population of ethnic group $k$ 888840 359480 820580 3048470 $N_{vk}$ Population of ethnic group $k$ with at least one vaccine (%) 58.0 76.5 76.8 84.0 $N_{vbk}$ Population of ethnic group $k$ with a booster (%) 20.9 30.9 40.1 51.7 cumulative reported cases per capita for ethnic group $k$ (%) 23.9 30.5 17.1 18.6
3.2.4 Proportionate mixing model estimation
We first considered the case where people mixed with all ethnicities in proportion to their population size. Ethnicity-specific attack rates (i.e., cumulative number of cases per capita) were calculated from the 1st of February to the 16th of May, 2022, to capture the first Omicron wave in Aotearoa. These attack rates were then scaled up using the assumed CAR. The four ethnicity-specific transmission rates, $a_{k}$, were chosen such that the attack rates from the model matched the attack rates from the case data by using the optimize.least_squares function from the SciPy package in Python with its default parameters. As there were 4 parameters and 4 attack rates to fit to, we found a unique solution for the transmission rates that provided an exact match to the target attack rates. All models were fit over a 365-day period, as it allowed sufficient time for the epidemic to run its course.
3.2.5 Assortative mixing model estimation
@@
We then considered the case where people were more likely to mix with others in the same ethnic group. A method to determine the 5th parameter, the assortativity constant, was needed. Statistical area data was used to select a value for the assortativity constant using a similar method as Ma et al. [ma_modeling_2021]. The method first constructed an empirically estimated contact matrix, $\boldsymbol{T}^{\prime}$, using a specified transmission vector, $\boldsymbol{a}$, as defined in the following equation:
$T_{ij}^{\prime}$ $=\sum_{l}\frac{a_{i}a_{j}N_{i,l}N_{j,l}}{\sum_{k}a_{k}N_{k,l}}$ (17)
where $N_{i,l}$ the population size of ethnicity group $i$ in statistical area $l$, and $T_{ij}^{\prime}$ is the total number of contacts between people in group $i$ and people in group $j$ (multiplied by the transmission probability).
@@
The definition of this matrix assumes that proportionate mixing occurs within each statistical area, and then aggregates the resulting contacts over the statistical areas. This was based on an approximating assumption that people mix proportionately with other people in the same geographical area and do not mix with people in other areas. This assumption was not entirely accurate to the real world but gave an approximate empirically based estimate of the average relative number of interactions an individual from a given ethnicity group would have.
When constructing this matrix, we assumed frequency-dependent interactions within statistical areas, whereas Ma et al. assumed density-dependent interactions in their block groups. Frequency-dependent interactions assume that the mean number of people an individual interacts with per day is independent of the population size in the census block or statistical area. In contrast, density-dependent interactions assume that the number of people an individual interacts with per day increases proportionally to the population size in the census block or statistical area in which they live. If the population blocks were all the same size, there would be no difference between these models, but when the blocks are different sizes, the density-based assumption would lead to an aggregated social contact matrix (after normalisation) that is disproportionately affected by the contact patterns within the larger blocks. The Aotearoa statistical areas have substantial variation in population size and, in this situation, we believe that the frequency-dependent paradigm is more realistic than the density-dependent one. This ignores systematic variations in contact rates with population density, for example between urban and rural areas, but these would also not be well captured by the density-dependent formulate of Ma et al. [ma_modeling_2021] as the most density populated statistical areas are not necessarily the ones with the largest population size (see Supplementary Figures 2 and 3).
@@
The values of $\boldsymbol{a}$ and $\epsilon$ were estimated jointly via the iterative method used by Me at al. [ma_modeling_2021] as follows. For a given transmission vector, $\boldsymbol{a}$, we found the assortativity value, $\epsilon$, that minimised the difference between the empirical matrix $T^{\prime}$ and the assortative mixing matrix $T$, defined by Equation 16.
@@
where $N_{i,l}$ is the population of ethnic group $i$ living in geographic region $l$. Then, for this value of epsilon, we chose the values of $\boldsymbol{a}$ for which the model matched with the target attack rates in the same way as described previously. This iterative process was repeated until the absolute difference between successive values of $\epsilon$ was less than $10^{-6}$.
3.2.6 Unconstrained mixing model estimation
Finally, we considered a model in which people were more likely to mix with others living in the same geographical area, but without imposing an assortative mixing model, which constrains all ethnicity groups to have the same assortativity, and assigns between-group interactions in proportion to population size. We refer to this as the unconstrained mixing model. To implement this model, we directly defined the transmission matrix in terms of the area-level population data via Equation 17. We then estimated the transmission rates as previously, by choosing the values of $a_{i}$ for which the model matched with the target attack rates.
Note that because the matrix in Equation 17 was formulated to have frequency-dependent contact rates, it has the required property that the the total average contact rate for an individual in group $i$ is equal to $a_{i}$, provided that $\sum_{l}N_{i,l}=N_{i}$ (see Supplementary Material section 3). This is not true of the density-dependent model used by Ma et al. [ma_modeling_2021], which can therefore only be used in conjunction with an assortative mixing assumption and not the unconstrained mixing model considered here. Because of the requirement that the area-level population data used to construct the matrix ($N_{i,l}$) aggregates to the national-level populations used in the model ($N_{i}$), we could only apply the unconstrained mixing model to the prioritised SA2 data and not the total ethnicity data.
@@ 3.3 Scenario analysis
We created four scenarios to quantify the effect that the ethnic disparity in vaccination rates, disparity in transmission rates, and the assortativity constant had on model results. These scenarios, seen in Table 3.3, took the assortative model fit and made one of the following changes: set $\epsilon=0$ to measure the effect of assortativity on the model; set all vaccination rates to the population-average vaccination rates; or set all transmission rates to the population-average transmission rates. These scenarios conserve the total number of vaccinations and the average transmission rate.
Code and data to replicate results can be found in the GitHub repository [lomas_modelling_2025].
@@ 4 Results
4.1 Proportionate mixing model
@@
As expected with an SEIR model, a single epidemic wave that ends as a result of herd immunity in the population was observed (Figure 2). Recall that the transmission rates were chosen to exactly match the attack rates observed in the data (after adjustment for the assumed CAR). As such, the final recovered population in each ethnicity matched the number of people infected according to the attack rate data and the assumed value for CAR (indicated by the dashed horizontal line).
@@
Figure 2: Per capita SEIR plots fit to data on confirmed cases under the assumption of proportionate mixing and an assumed CAR of 50% for a) Māori, b) Pacific, c) Asian, and d) Europeans/Other. The dashed line is the cumulative attack rate according to the data on reported cases and the assumed CAR of 50%.
@@
For all the assumed values of CAR, we chose transmission rates to exactly match the inferred attack rate values; see Figure 3 and Table 4.3. For all values of CAR investigated, the Pacific transmission rate was highest, followed by the Māori, European/Other, and then Asian rates, respectively. The initial reproduction number increased with a decrease in the assumed CAR and was highest when Māori and Pacific Peoples were given a higher CAR than the other groups.
4.2 Assortative mixing model